When a man observed a sobriety checkpoint conducted by a police department, he saw 654 drivers were screened and 8 were arrested for driving while intoxicated. Based on those results, we can estimate that P(w) 0.01223, where W denotes the event of screening a driver and getting someone who is intoxicated. What does P (W) denote, and what is its value? What does P() represent? OA. P(W) denotes the probability of screening a driver and finding that he or she is intoxicated. B. P(W) denotes the probability of screening a driver and finding that he or she is not intoxicated. O c. P (W) denotes the probability of a driver passing through the sobriety check point. D. P(W) denotes the probability of driver being intoxicated. P(W)- (Round to five decimal places as needed.)
Answer #1

P ( \bar{w} ) = Complement of P ( w )

= > Which is the probability of screening a driver and finding that he or she is not intoxicated.

P ( \bar{w} ) = 1 - P( w )

= 1 - 0.01223

= 0.98777

P ( \bar{w} ) = ( 654 - 8 ) / 654 =   0.98777